3. For any piecewise continuous function f that is p-periodic, verify that a+p р Jf(x)dx 0 а 2

please i need the question 3 for the detailed  proof ! Thanks!
Answer #1

37 fis period having t is periodic function period p. = f(u) = f(a+b) Stude atb f(a+x) dx f(n) da T = f(u) da atp-a-& I f(ata